OpenStudy (anonymous):

if sin0=4/5 what is the value of cos0 in quadrant1?

OpenStudy (anonymous):

3/5

OpenStudy (anonymous):

if cos0=5/13 what is the value of sin0 in quadrant1?

OpenStudy (anonymous):

12/13

OpenStudy (haseeb96):

Please explain it step by step

OpenStudy (anonymous):

id sin0=15/17 and cos 0=8/17 find the value of tan0 in quadrant 1?

OpenStudy (anonymous):

explain that one step by step please

OpenStudy (haseeb96):

tan0=15\8

OpenStudy (haseeb96):